OpenStudy (anonymous):

The slope of the straight line passing through the points (0,0) and (3,2) is: 2/3

OpenStudy (kenljw):

Yes, the slope is the change in y divided by the change in x, the y intercept is when x =0 which is y = 0

OpenStudy (anonymous):

If you are asking this in engineering, there are two logical possibilities. A) you are not an engineering student B) you are really naive/dumb.

OpenStudy (anonymous):

in this case x=3-0=3 y=2-0=2 slope \[\tan^{-1} \ y/x ]